For gee-whiz devices calculate the following: working capital, current ratio, and acid-test ratio.

2-29 For Gee-Whiz Devices calculate the following working capital, current ratio, and acid-test ratio Gee-Whiz Devices Balance Sheet Data Cash Market securities Net accounts and notes receivable Retailers inventories Prepaid expenses Accounts and notes payable (short-term) Accrued expenses $100,000 45,000 150,000 200,000 8,000 315,000 90,000
Answer #1

1) Working Capital = Current Asset - Current Liabilities

Current Asset = Cash+Marketable Securities+Net accounts and notes receivable+Retailers Inventories+Prepaid expences

= 100000+45000+150000+200000+8000

= $ 503000 -------------------------------------------------------------------(1)

Current Liabilities = Accounts and notes payable(Short term) + Accrued expenses

= 315000+90000

= $ 405000-------------------------------------------------------------------(2)

Working Capital = Current Asset - Current Liabilities

503000 - 405000

= $ 98000

2) Current Asset ratio = Current Asset/Current Liabilities

As we already calculated CA and CL above in eqation 1 and 2

Current Asset ratio = Current Asset/Current Liabilities

Current Asset ratio = 503000/405000

= 1.24:1

3) Acid test Ratio = Quick Asset/Quick Liabilities

Quick Asset = Current Asset - ( Stock+ Prepaid expences)

= 503000- (200000+8000)

= $ 295000

Quick Liabilities = Current Liabilities - (Bank Overdraft+Cash credit)

= 405000 - (0+0)

= $ 405000

Acid test Ratio = Quick Asset/Quick Liabilities

= 295000/405000

= 0.728.